You are long 10 Sep mini S&P 500 futures contracts at a cost of 2825. The...

  1. You are long 10 Sep mini S&P 500 futures contracts at a cost of 2825. The price is now 2999.75. What is your unrealized gain or loss?
  1. You are going to short 1 Dec US Treasury Bond futures contract. The price quote is 156-13. What is that in decimals? What is the value of this 1 contract in dollars?
  1. You heard that it only costs about .864 euros to buy one dollar. At approximately what price would you expect the Sep Euro contract to be trading?
  1. If you thought that the Euro was going to rise relative to the British Pound, how would you design a trade using the contracts above?
Answer #1

The long position is in profit if the futures contract increases in price

The S&P 500 mini futures contract has a contract multiplier of 50. That is, for each $1 change in S&P500 futures price, the gain/loss is $50

Profit = increase in futures price * number of contracts * multiplier = (2999.75 - 2825) * 10 * 50 = 87375

Treasury Bond futures are quoted in fractions of 32. In this case, the quote of 156-13 means that the quote is 156 plus 13/32.

Quote in decimals = 156 + (13 / 32) = 156.40625

The face value of Treasury Bond futures is $100,000. Value of this contract = quote * face value = $156.40625 * $100,000 = $15,640,625

The Sep Euro contract would trade slightly higher or slightly lower than 0.864. It would not be exactly equal to the spot price due to differences in inflation, interest rates, transaction costs etc.

If the Euro is going to rise relative to the British Pound, the trade to make a profit from such a rise would be :

  • Buy the Euro/USD futures contract
  • Sell the GBPUSD futures contract